OpenStudy (anonymous):

Use the given function value(s) and the trigonometric identities to find the indicated trigonometric functions. tan β = 5 (a) cot β (b) cos β (c) tan(90°-β) (d) csc β

OpenStudy (anonymous):



OpenStudy (anonymous):

cot B = 1/5

OpenStudy (anonymous):

cos B = 1/sqrt(26)

OpenStudy (anonymous):

(b) i got sq rt(26)/26

OpenStudy (anonymous):

csc B = sqrt(26)/5

OpenStudy (anonymous):

tan (90 - B) = 1/5

OpenStudy (anonymous):

sqrt(26)/26 = 1/sqrt(26)

OpenStudy (anonymous):

oh so its the same?

OpenStudy (anonymous):

Never mind i understand now. thanks!

OpenStudy (anonymous):

they are the same..if you rationalize the denominator of 1/sqrt(26) you get sqrt(26)/26

OpenStudy (anonymous):

Oh by the way for (d) i got 5 sqrt (26)/ 26.